# Possible Liquid-Nitrogen-Temperature Superconductivity Driven by Perpendicular Electric Field in the Single-Bilayer Film of La3Ni2O7 at Ambient Pressure

### v2

@@ V Code availability
VI Acknowledgement
We are grateful to the stimulating discussions with Chen Lu. F. Y. and C. W are supported by the National Natural Science Foundation of China (NSFC) under the Grant No. 12234016, and also supported by the NSFC under the Grant Nos. 12074031 and 12174317, respectively. D. X. Y. is supported by NKRDPC-2022YFA1402802, NSFC-92165204, and Guangdong Provincial Quantum Science Strategic Initiative (GDZX2401010).
VII Author contributions
F. Yang proposed the main idea and supervised the study. Z.-Y. Shao performed the SBMF study. J.-H. Ji performed the DMRG study. F. Yang, Z.-Y. Shao and J.-H. Ji wrote the paper. All authors significantly contributed to the data analysis and discussion. .

### v3

@@ Appendix F Code availability
Appendix G Acknowledgement
We are grateful to the stimulating discussions with Chen Lu. F. Y. and C. W. is supported by the National Natural Science Foundation of China (NSFC) under the Grant No. 12234016. F. Y. is also supported by the CAS Superconducting Research Project under Grant No. [SCZX-0101] and the NSFC under the Grant No. 12074031. C. W. is also supported by the NSFC under the Grant No. 12174317. D. X. Y. is supported by NSFC-12494591, NSFC-92165204, NSFC-92565303, NKRDPC-2022YFA1402802, Research Center for Magnetoelectric Physics of Guangdong Province (2024B0303390001), and Guangdong Provincial Quantum Science Strategic Initiative (GDZX2401010). C. W. is also supported by the New Cornerstone Science Foundation.
Appendix H Author contributions
F. Yang proposed the main idea and supervised the study. Z.-Y. Shao performed the SBMF study. J.-H. Ji performed the DMRG study. D.-X. Yao and C. Wu helped shape the main idea. F. Yang, Z.-Y. Shao and J.-H. Ji wrote the paper.
Appendix I Additional Information
Competing Interests
The authors declare no competing interests.